OpenStudy (anonymous):

The Perimeter Of A Rectangle Is Twice The The Sum Of Its Length And Its Width. The Perimeter Is 46 Meters And Its Length Is 2 Meters More Than Twice Its Width. What Is The Length ?

OpenStudy (anonymous):

16 I Dont Need Help

Parth (parthkohli):

\[\rm \begin{aligned} l &= 2 + w \\ \rm l + w & = 46\end{aligned}\]

Parth (parthkohli):

16 is the wrong answer

OpenStudy (anonymous):

No I Got It Correct

Parth (parthkohli):

Nope.

Parth (parthkohli):

2(16 + 14) = 60

OpenStudy (anonymous):

Nope I Got The Right Answer On The Computer

Parth (parthkohli):

Nope.

OpenStudy (anonymous):

I Fucccin Got It Right Okay

OpenStudy (anonymous):

Lol, someone is ignorant, Parth knows what he is doing lady

OpenStudy (anonymous):

Hold On Ima Show Yall

OpenStudy (anonymous):

sounds good

Parth (parthkohli):

I just proved you that your solution was wrong. *sigh*

OpenStudy (anonymous):

I have a strong urge to make a pun on her user name :P

OpenStudy (anonymous):

Umm Both Of Youu Btttchs Can Fuccin Move Yall Mad Bcuss Im Right Nd Even If It Was Wrong R Yalll Mad

Parth (parthkohli):

So stubborn...

OpenStudy (anonymous):

somebody just went all gansta in a math help website

OpenStudy (anonymous):

yu so bad gall

OpenStudy (aripotta):

i don't think anybody can understand what you're saying :/

Parth (parthkohli):

Solve this and tell me what you get:\[2(x + x - 2) = 46\]

OpenStudy (anonymous):

and yuu lil biscut... you are such a female dog that you cant see he is right

OpenStudy (anonymous):

46=2w+4w+4 6w=46-4 6w=42 w=7 L=2*7+2 L=16 So How You Gone Tell Me

OpenStudy (anonymous):

I am gona tell you just like :"you are WRONG!" ..... thats how :]

Parth (parthkohli):

LOOOOOOOOOOOOOOOOOOOOOOOLLLLLLLLLL OMG

Parth (parthkohli):

Look at the question, and your answer.

OpenStudy (anonymous):

Lhhh Right Thanks So R You Mad And @ParthKohli Its 25/2

OpenStudy (anonymous):

Okay If Your So Right Whats The Answer

Parth (parthkohli):

Nvm, I was just trolling. :D

OpenStudy (anonymous):

this amuses me greatly...

OpenStudy (anonymous):

I am pleased by your agony!

OpenStudy (anonymous):

I Thought So You All Are DISMISSED

Parth (parthkohli):

@DamnImBad I was just trying to test you; it's not 25/2.

Parth (parthkohli):

Yeah, it's 16. :P

OpenStudy (aripotta):

no it's not?

OpenStudy (anonymous):

*white boy shuffle* VICTORY!

Parth (parthkohli):

It is.

OpenStudy (anonymous):

But It Really Is

Parth (parthkohli):

2(16 + 7) = 46

OpenStudy (aripotta):

let me check

Parth (parthkohli):

You have passed. \(\checkmark\)

OpenStudy (anonymous):

TF Im Not Dumb So Please Move

OpenStudy (anonymous):

are u sure?

OpenStudy (aripotta):

oh yea it is lol. why'd you guys do that to her?

OpenStudy (anonymous):

you might be and not even know it!

Parth (parthkohli):

Yes, it's 16.

Parth (parthkohli):

Next test:

Parth (parthkohli):

What was the fallacy in my reply earlier?

OpenStudy (anonymous):

PLEASE GET OFF

OpenStudy (anonymous):

what is the air velocity of a English sparrow?

OpenStudy (anonymous):

WHATS YOUR FAVORITE COLOR!

Parth (parthkohli):

Hehe nvm sorry for the trouble

OpenStudy (anonymous):

Tf What R You Talking About @DrakeS

OpenStudy (anonymous):

I am quoting a movie... apparently none of you have seen *face palm*

OpenStudy (anonymous):

No So Move

OpenStudy (anonymous):

this proves you are stupid girl...

OpenStudy (anonymous):

you dont know mony python!

OpenStudy (anonymous):

wall, what do ya know. She is speechless!

OpenStudy (anonymous):

Btttch Move Tf GooodBye Why R You The Only One Still Here

OpenStudy (anonymous):

Because I enjoy talking to you, why else?

OpenStudy (anonymous):

Tf Who R You

OpenStudy (anonymous):

now... if only we could fix your grammar issues

OpenStudy (anonymous):

Now Only If Youu Didnn Want Yo retrice Whooped Youu Better Move Nd Leave Me The Fucc Alone

OpenStudy (anonymous):

or wut? U gonna insult me with bad grammar?

OpenStudy (anonymous):

Btttch Ima Talk Bad Tf Youu Dnn Knw Me Btttch I Will Spit IN Yo Face Puuusyy Youu Need Shake The Mutha Fucccin Spot

OpenStudy (anonymous):

oh snap, I only understood half that and it was enough to get all scared! Ima back off female dog, b4 u get all gansta on me booty!

OpenStudy (anonymous):

Yeaaa

OpenStudy (anonymous):

I lied, this is too much fun to stop now!

OpenStudy (anonymous):

btw, how did u get so ugly?

OpenStudy (anonymous):

oh, im sorry, I used too large words for you to understand...

OpenStudy (anonymous):

hw u so ugly?

OpenStudy (anonymous):

then she mails me "Lhhh Btttch I Will Spit In Ya Face Nd Kill Yo Mama "

OpenStudy (anonymous):

then "Okay Btttch Youu Think Im Playinn "

OpenStudy (anonymous):

and the conversation continues in mail :P

OpenStudy (anonymous):

LET ME LOVE YOU DRAKE

OpenStudy (anonymous):

U hittin on me woman?

OpenStudy (anonymous):

well, sorta. maybe figuratively :D

OpenStudy (anonymous):

Get Tf Off My pellett

OpenStudy (anonymous):

My Shhittt

OpenStudy (anonymous):

lol, leave yo name and numba and I will contact you as soon as I become indian

OpenStudy (anonymous):

jk, i already have a gf

OpenStudy (anonymous):

My dear girl. if you talk by capitalizing every damn word you type, and talk pellet like tha,t and then ask pellet questions like that, and then fluttering answer the question yourself, you will get trolled by every creative person on this site. DEAL WITH IT.

OpenStudy (anonymous):

hehe, so true

OpenStudy (anonymous):

and then, if you don't calm your tits fast enough, people are gonna MAKE SURE you're chucked out of OS. I will personally oversee it.

OpenStudy (anonymous):

Lhhh Bttttch Wasssuhp Ima Tell Youu Like I Told Him Btttch I Will Spit N Yo Face Killl Yo Mamam Nd Yoo Familly Btttch Youuu Better Move TF

OpenStudy (anonymous):

me too, i have been for a few minuets now :P

OpenStudy (anonymous):

haha, you sure can type fast tanv

OpenStudy (anonymous):

mofo I DOUBLE DARE you to do bull* like this again. you better stop typing. I know the admin dudes on this site personally, i'll get you chucked out.

OpenStudy (anonymous):

haha, I got this,

OpenStudy (anonymous):

So Fucccin Move Off My pellettttttttttttt

OpenStudy (anonymous):

keep typin female dog!

OpenStudy (anonymous):

Shhiiitt

OpenStudy (anonymous):

yaaaa, tahts right! I WIN

OpenStudy (anonymous):

Drake, I don't get it. whose side are you on??? :P

OpenStudy (anonymous):

Lhhh She Knww Themm So Personally Bcuss She Always Givvvin BLOW JOBS BTtttch Byee

OpenStudy (anonymous):

I am on whatever side gets her the most mad :P

OpenStudy (anonymous):

oh snap, u so classy woman!

OpenStudy (anonymous):

cant even insult people properly :P